# Sample size and power calculations for causal inference in observational studies

### v3

@@ A.4 Proof of Proposition 2
$\mathbb{E}\left[\frac{\{Y(0)-\mathbb{E}[Y(0)]\}^{2}}{1-e(X)}\right]=\mathbb{E}\left[\frac{a_{0}^{2}(W_{e}-\mu_{e})^{2}+\epsilon_{0}^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right]=a_{0}^{2}\mathbb{E}\left[\frac{(W_{e}-\mu_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{0}^{2}\mathbb{E}\left[\frac{1}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right],$ (A.33)
where
$\mathbb{E}\left[\frac{1}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ $=\int\left(1+e^{w}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(w-\mu_{e})^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}w$ (A.34)
$=\int\left(1+e^{u+\mu_{e}}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u$ (A.35)
$=\int\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u+\int\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(u-\sigma_{e}^{2})^{2}}{2\sigma_{e}^{2}}+\mu_{e}+\frac{\sigma_{e}^{2}}{2}\right)\,\mathrm{d}u$ (A.36)
$=1+\exp\left(\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right),$ (A.37)
and
$\mathbb{E}\left[\frac{(W_{e}-\mu_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ $=\int(w-\mu_{e})^{2}\left(1+e^{w}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(w-\mu_{e})^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}w$ (A.38)
$=\int u^{2}\left(1+e^{u+\mu_{e}}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u$ (A.39)
$=\int u^{2}\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u+\int u^{2}\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(u-\sigma_{e}^{2})^{2}}{2\sigma_{e}^{2}}+\mu_{e}+\frac{\sigma_{e}^{2}}{2}\right)\,\mathrm{d}u$ (A.40)
$=\sigma_{e}^{2}+(\sigma_{e}^{4}+\sigma_{e}^{2})\exp\left(\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right).$ (A.41)
Therefore,
$\mathbb{E}\left[\frac{\{Y(0)-\mathbb{E}[Y(0)]\}^{2}}{1-e(X)}\right]=a_{0}^{2}\sigma_{e}^{2}+\sigma_{0}^{2}+\left[a_{0}^{2}\sigma_{e}^{2}(\sigma_{e}^{2}+1)+\sigma_{0}^{2}\right]\exp\left(\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right).$ (A.42)
Adding (A.32) and (A.42), we have
$\small V=(a_{1}^{2}+a_{0}^{2})\sigma_{e}^{2}+(\sigma_{1}^{2}+\sigma_{0}^{2})+\left[a_{1}^{2}\sigma_{e}^{2}(\sigma_{e}^{2}+1)+\sigma_{1}^{2}\right]\exp\left(-\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right)+\left[a_{0}^{2}\sigma_{e}^{2}(\sigma_{e}^{2}+1)+\sigma_{0}^{2}\right]\exp\left(\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right).$ (A.43)
A.7 Proof that the calculated sample size increases with $\rho^{2}$.
We rewrite $\widetilde{V}=\widetilde{V}_{\mathrm{SH}}+\rho^{2}\widetilde{V}_{\mathrm{adj}}$, where $\widetilde{V}_{\mathrm{SH}}=2+\exp(-\mu_{e}+\sigma_{e}^{2}/2)+\exp(\mu_{e}+\sigma_{e}^{2}/2)$ and
$\widetilde{V}_{\mathrm{adj}}=\frac{\sigma_{e}^{2}}{\sigma_{e\mid 1}^{2}}+\frac{\sigma_{e}^{2}}{\sigma_{e\mid 0}^{2}}-2+\left[\frac{\sigma_{e}^{2}(\sigma_{e}^{2}+1)}{\sigma_{e\mid 1}^{2}}-1\right]\exp\left(-\mu_{e}+\frac{\sigma_{e}^{2}}{2}\right)+\left[\frac{\sigma_{e}^{2}(\sigma_{e}^{2}+1)}{\sigma_{e\mid 0}^{2}}-1\right]\exp\left(\mu_{e}+\frac{\sigma_{e}^{2}}{2}\right).$ (A.44)
Recall that $W\sim\mathsf{N}(\mu_{e},\sigma_{e}^{2})$. Then, $\sigma_{e\mid z}^{2}=\mathbb{V}[W\mid Z=z]\leq\mathbb{V}[W]=\sigma_{e}^{2}$. Hence,
$\frac{\sigma_{e}^{2}}{\sigma_{e\mid z}^{2}}\geq 1,\quad\frac{\sigma_{e}^{2}(\sigma_{e}^{2}+1)}{\sigma_{e\mid z}^{2}}-1\geq\frac{\sigma_{e}^{2}}{\sigma_{e\mid z}^{2}}-1\geq 0.$ (A.45)
Therefore, $\widetilde{V}_{\mathrm{adj}}\geq 0$.
A.8 Proof of the asymptotic variance of the general Hájek WATE estimator.
@@
$\phi=\begin{pmatrix}\phi_{1}\\ \phi_{0}\\ \psi_{\alpha}\end{pmatrix}\quad\text{with}\quad\begin{aligned} \phi_{1}&=Z_{i}\cdot w_{1}(X_{i};\alpha)\cdot\{Y_{i}(1)-\xi_{1}\},\\ \phi_{0}&=(1-Z_{i})\cdot w_{0}(X_{i};\alpha)\cdot\{Y_{i}(0)-\xi_{0}\},\\ \psi_{\alpha}&=e^{\prime}(X_{i};\alpha)\{Z_{i}-e(X_{i};\alpha)\}.\end{aligned}$ (A.46)
@@
$\sqrt{N}\begin{pmatrix}\hat{\xi}_{1}^{-}\xi_{1}^{*}\\ \hat{\xi}_{0}^{-}\xi_{0}^{*}\\ \hat{\alpha}-\alpha^{*}\end{pmatrix}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,(A^{*})^{-1}B^{*}(A^{*})^{-\top}),$ (A.47)
where $A^{*}$ and $B^{*}$ are the values of $-\mathbb{E}\left(\frac{\partial\phi}{\partial(\xi_{1},\xi_{0},\alpha)}\right)$ and $\mathbb{E}[\phi\phi^{\top}]$ evaluated at $(\xi_{1}^{*},\xi_{0}^{*},\alpha^{*})$ respectively.
@@
$\frac{\partial\phi}{\partial(\xi_{1},\xi_{0},\alpha)}\big{|}_{(\xi_{1}^{*},\xi_{0}^{*},\alpha^{*})}=\begin{pmatrix}\frac{\partial\phi_{1}^{*}}{\partial\xi_{1}}&0&\frac{\partial\phi_{1}^{*}}{\partial\alpha^{\top}}\\ 0&\frac{\partial\phi_{0}^{*}}{\partial\xi_{0}}&\frac{\partial\phi_{0}^{*}}{\partial\alpha^{\top}}\\ 0&0&\frac{\partial\psi_{\alpha}^{*}}{\partial\alpha^{\top}},\end{pmatrix}$ (A.48)
@@
where
$a_{11}$ $=\mathbb{E}\left[Z_{i}\cdot w_{1}(X_{i};\alpha^{*})\right]=\mathbb{E}[e(X_{i};\alpha^{*})w_{1}(X_{i};\alpha^{*})]=\mathbb{E}[h(X_{i};\alpha^{*})],$
$a_{22}$ $=\mathbb{E}\left[(1-Z_{i})\cdot{w_{0}(X_{i};\alpha^{*})}\right]=\mathbb{E}[(1-e(X_{i};\alpha^{*}))w_{0}(X_{i};\alpha^{*})]=\mathbb{E}[h(X_{i};\alpha^{*})],$
$a_{13}$ $=\mathbb{E}\left[-Z_{i}\cdot\{Y_{i}(1)-\xi_{1}^{*}\}\cdot w_{1}^{\prime}(X_{i};\alpha^{*})\right]=-\mathbb{E}\left[e(X_{i};\alpha^{*})\cdot\{Y_{i}(1)-\xi_{1}^{*}\}\cdot w_{1}^{\prime}(X_{i};\alpha^{*})\right],$
$a_{23}$ $=\mathbb{E}\left[-(1-Z_{i})\cdot\{Y_{i}(0)-\xi_{0}^{*}\}\cdot w_{0}^{\prime}(X_{i};\alpha^{*})\right]=-\mathbb{E}\left[(1-e(X_{i};\alpha^{*}))\cdot\{Y_{i}(0)-\xi_{0}^{*}\}\cdot w_{0}^{\prime}(X_{i};\alpha^{*})\right],$
$a_{33}$ $=\mathbb{E}\left[-e^{\prime}(X_{i};\alpha^{*})e^{\prime}(X_{i};\alpha^{*})^{\top}\right].$
@@
$b_{22}$ $=\mathbb{E}\left[(1-Z_{i})\cdot{\{Y_{i}(0)-\xi_{0}^{*}\}^{2}}\cdot{w_{0}(X_{i};\alpha^{*})^{2}}\right]=\mathbb{E}\left[{\{Y_{i}(0)-\xi_{0}^{*}\}^{2}}\{1-e(X_{i};\alpha^{*})\}{w_{0}(X_{i};\alpha^{*})}^{2}\right],$
$b_{13}$ $=\mathbb{E}\left[Z_{i}(Z_{i}-e(X_{i};\alpha^{*}))\cdot w_{1}(X_{i};\alpha^{*})\cdot\{Y_{i}(1)-\xi_{1}^{*}\}\cdot e^{\prime}(X_{i};\alpha^{*})\right]$
$=\mathbb{E}\left[e(X_{i};\alpha^{*})\{1-e(X_{i};\alpha^{*})\}\cdot w_{1}(X_{i};\alpha^{*})\cdot\{Y_{i}(1)-\xi_{1}^{*}\}\cdot e^{\prime}(X_{i};\alpha^{*})\right],$
$b_{23}$ $=\mathbb{E}\left[(1-Z_{i})(Z_{i}-e(X_{i};\alpha))\cdot w_{0}(X_{i};\alpha^{*})\cdot\{Y_{i}(0)-\xi_{0}^{*}\}\cdot e^{\prime}(X_{i};\alpha^{*})\right]$
$=\mathbb{E}\left[-e(X_{i};\alpha^{*})\{1-e(X_{i};\alpha^{*})\}\cdot w_{0}(X_{i};\alpha^{*})\cdot\{Y_{i}(0)-\xi_{0}^{*}\}\cdot e^{\prime}(X_{i};\alpha^{*})\right],$
$b_{33}$ $=\mathbb{E}\left[e(X_{i};\alpha^{*})\{1-e(X_{i};\alpha^{*})\}e^{\prime}(X_{i};\alpha^{*})e^{\prime}(X_{i};\alpha^{*})^{\top}\right].$
@@
$(A^{*})^{-1}B^{*}(A^{*})^{-\top}$ $=\begin{pmatrix}a_{11}^{-1}&0&-a_{11}^{-1}a_{13}a_{33}^{-1}\\ 0&a_{11}^{-1}&-a_{11}^{-1}a_{23}a_{33}^{-1}\\ 0&0&a_{33}^{-1}\end{pmatrix}\begin{pmatrix}b_{11}&0&b_{13}\\ 0&b_{22}&b_{23}\\ b_{13}^{\top}&b_{23}^{\top}&b_{33}\end{pmatrix}\begin{pmatrix}a_{11}^{-1}&0&0\\ 0&a_{11}^{-1}&0\\ -a_{33}^{-1}c_{13}^{\top}a_{11}^{-1}&-a_{33}^{-1}c_{23}^{\top}a_{11}^{-1}&a_{33}^{-1}\end{pmatrix}$
@@
A.9 Proof of (C.5)
Recall that $Y(z)=a_{z}W_{e}+\epsilon_{z}$, where $\epsilon_{z}\sim\mathsf{N}(\mu_{z},\sigma_{z}^{2})$. Then,
$\xi_{z}=\frac{\mathbb{E}[h(X)Y(z)]}{\mathbb{E}[h(X)]}=a_{z}\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}+\frac{\mathbb{E}[h(W_{e})\epsilon_{z}]}{\mathbb{E}[h(W_{e})]}=a_{z}\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}+\mu_{z}.$ (A.49)
The last equation holds because $\epsilon_{z}$ and $W_{e}$ are uncorrelated normal random variables and thus independent. Then,
$\mathbb{E}\left[\frac{(Y_{i}(1)-\xi_{1})^{2}}{e(X_{i})}h(X_{i})^{2}\right]$ $=\mathbb{E}\left[\left\{a_{1}\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)+\epsilon_{1}-\mu_{1}\right\}^{2}\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ (A.50)
$=\mathbb{E}\left[\mathbb{E}\left[\left\{a_{1}\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)+\epsilon_{1}-\mu_{1}\right\}^{2}\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\mid W_{e}\right]\right]$ (A.51)
$=\mathbb{E}\left[\left\{a_{1}^{2}\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)^{2}+\sigma_{1}^{2}\right\}\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ (A.52)
$=a_{1}^{2}\mathbb{E}\left[\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)^{2}\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{1}^{2}\mathbb{E}\left[\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right].$ (A.53)
Similarly,
$\mathbb{E}\left[\frac{(Y_{i}(0)-\xi_{0})^{2}}{1-e(X_{i})}h(X_{i})^{2}\right]$ $=a_{0}^{2}\mathbb{E}\left[\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)^{2}\frac{h(W_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{0}^{2}\mathbb{E}\left[\frac{h(W_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right].$ (A.54)
B Examples
B.1 Normality of linear combination of covariates.
We consider the simulated data in Section 4 and three real world datasets: (i) RHC; (ii) the Best Apnea Interventions for Research (BestAIR) trial, a individually randomized trial with 169 units, with 9 continuous or binary covariates; more details can be found in Zeng et al., (2021); (iii) HSR: an observational study on the racial disparities of breast cancer screening, with 56,480 units and 17 binary covariates; more details can be found in Li et al., (2013). For each dataset, we fit a simple logistic model to estimate the propensity score $e(X)=\operatorname{\mathrm{e}xpit}(\beta_{0}+X^{\prime}\beta)$, and we provide the density plot and Q-Q plot in Figure 4.
Figure 4: Density plot and Q-Q plot of fitted logit propensity scores, which is a linear combination of covariates. The proximity of the points to the reference line in the Q-Q plot shows the closeness of the distribution to a normal distribution.

### v4

@@ A.7 Proof of the analytical form of variance $V$ in Theorem 2
$\mathbb{E}\left[\frac{\{Y(0)-\mathbb{E}[Y(0)]\}^{2}}{1-e(X)}\right]=a^{2}\sigma_{e}^{2}+\sigma_{y}^{2}+\left[a^{2}\sigma_{e}^{2}(\sigma_{e}^{2}+1)+\sigma_{y}^{2}\right]\exp\left(\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right).$ (64)
Adding (63) and (64), we have
$V=2(a^{2}\sigma_{e}^{2}+\sigma_{y}^{2})+2\left[a^{2}\sigma_{e}^{2}(\sigma_{e}^{2}+1)+\sigma_{y}^{2}\right]\exp\left(\frac{\sigma_{e}^{2}}{2}\right)\cosh(\mu_{e}).$ (65)
Plugging in the expression for $a$ and $\sigma_{y}^{2}$,
$V=2S^{2}\left\{1+(\rho^{2}\sigma_{e}^{2}+1)\exp\left(\frac{\sigma_{e}^{2}}{2}\right)\cosh(\mu_{e})\right\}.$ (66)
A.8 Proof of the asymptotic variance of the general Hájek WATE estimator
@@
$\phi=\begin{pmatrix}\phi_{1}\\ \phi_{0}\\ \psi_{\alpha}\end{pmatrix}\quad\text{with}\quad\begin{aligned} \phi_{1}&=Z_{i}\cdot w_{1}(X_{i};\alpha)\cdot\{Y_{i}(1)-\xi_{1}\},\\ \phi_{0}&=(1-Z_{i})\cdot w_{0}(X_{i};\alpha)\cdot\{Y_{i}(0)-\xi_{0}\},\\ \psi_{\alpha}&=e^{\prime}(X_{i};\alpha)\{Z_{i}-e(X_{i};\alpha)\}.\end{aligned}$ (67)
@@
$\sqrt{N}\begin{pmatrix}\hat{\xi}_{1}^{-}\xi_{1}^{*}\\ \hat{\xi}_{0}^{-}\xi_{0}^{*}\\ \hat{\alpha}-\alpha^{*}\end{pmatrix}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,(A^{*})^{-1}B^{*}(A^{*})^{-\top}),$ (68)
where $A^{*}$ and $B^{*}$ are the values of $\mathbb{E}\left(\frac{\partial\phi}{\partial(\xi_{1},\xi_{0},\alpha)}\right)$ and $\mathbb{E}[\phi\phi^{\top}]$ evaluated at $(\xi_{1}^{*},\xi_{0}^{*},\alpha^{*})$ respectively.
@@
$\frac{\partial\phi}{\partial(\xi_{1},\xi_{0},\alpha)}\big|_{(\xi_{1}^{*},\xi_{0}^{*},\alpha^{*})}=\begin{pmatrix}\frac{\partial\phi_{1}^{*}}{\partial\xi_{1}}&0&\frac{\partial\phi_{1}^{*}}{\partial\alpha^{\top}}\\ 0&\frac{\partial\phi_{0}^{*}}{\partial\xi_{0}}&\frac{\partial\phi_{0}^{*}}{\partial\alpha^{\top}}\\ 0&0&\frac{\partial\psi_{\alpha}^{*}}{\partial\alpha^{\top}},\end{pmatrix}$ (69)
@@
where $a_{11}=\mathbb{E}\left[Z_{i}\cdot w_{1}(X_{i};\alpha^{*})\right]=\mathbb{E}[e(X_{i};\alpha^{*})w_{1}(X_{i};\alpha^{*})]=\mathbb{E}[h(X_{i};\alpha^{*})].$
@@
$b_{22}$ $=\mathbb{E}\left[(1-Z_{i})\cdot{\{Y_{i}(0)-\xi_{0}^{*}\}^{2}}\cdot{w_{0}(X_{i};\alpha^{*})^{2}}\right]$
$=\mathbb{E}\left[{\{Y_{i}(0)-\xi_{0}^{*}\}^{2}}\{1-e(X_{i};\alpha^{*})\}{w_{0}(X_{i};\alpha^{*})}^{2}\right].$
@@
$(A^{*})^{-1}B^{*}(A^{*})^{-\top}$
$=$ $\begin{pmatrix}a_{11}^{-1}&0&-a_{11}^{-1}a_{13}a_{33}^{-1}\\ 0&a_{11}^{-1}&-a_{11}^{-1}a_{23}a_{33}^{-1}\\ 0&0&a_{33}^{-1}\end{pmatrix}\begin{pmatrix}b_{11}&0&b_{13}\\ 0&b_{22}&b_{23}\\ b_{13}^{\top}&b_{23}^{\top}&b_{33}\end{pmatrix}\begin{pmatrix}a_{11}^{-1}&0&0\\ 0&a_{11}^{-1}&0\\ -a_{33}^{-1}c_{13}^{\top}a_{11}^{-1}&-a_{33}^{-1}c_{23}^{\top}a_{11}^{-1}&a_{33}^{-1}\end{pmatrix}$
@@
A.9 Proof of (19)
See Appendix 12 that $Y(z)=aW_{e}+z\tau+\epsilon$, where $\epsilon=W_{e}^{\perp}+\epsilon_{ori}$ is the compound error with $\operatorname{\mathbb{E}}[\epsilon]=\mu$ and $\operatorname{\mathbb{V}}[\epsilon]=\sigma_{y}^{2}$. $\epsilon_{ori}$ is the error term defined in Model (11). Then,
$\xi_{z}=\frac{\mathbb{E}[h(X)Y(z)]}{\mathbb{E}[h(X)]}$ $=a\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}+\frac{\mathbb{E}[h(W_{e})z\tau]}{\mathbb{E}[h(W_{e})]}+\frac{\mathbb{E}[h(W_{e})\epsilon]}{\mathbb{E}[h(W_{e})]}$ (70)
$=a\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}+z\tau+\mu.$ (71)
The last equation holds because of Lemma 4 (2). Then,
$\mathbb{E}\left[\frac{(Y_{i}(1)-\xi_{1})^{2}}{e(X_{i})}h(X_{i})^{2}\right]$
$=$ $\mathbb{E}\left[\left\{a\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)+\epsilon-\mu\right\}^{2}\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ (72)
$=$ $\mathbb{E}\left[\mathbb{E}\left[\left\{a\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)+\epsilon-\mu\right\}^{2}\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\mid W_{e}\right]\right]$ (73)
$=$ $\mathbb{E}\left[\left\{a^{2}\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)^{2}+\sigma_{y}^{2}\right\}\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ (74)
$=$ $a^{2}\mathbb{E}\left[\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)^{2}\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{y}^{2}\mathbb{E}\left[\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right].$ (75)
The third equation holds because $\operatorname{\mathbb{V}}[\epsilon|W_{e}]=\operatorname{\mathbb{V}}[\epsilon]$ and the cross term equals zero, see Appendix A.7. Similar derivation yields the other equation on $Y_{i}(0)$.
Appendix B Examples
B.1 Empirical examples of normality of linear combination of covariates
We consider the simulated data in Section 5 and three real world datasets: (i) RHC; (ii) the Best Apnea Interventions for Research (BestAIR) trial, a individually randomized trial with 169 units, with 9 continuous or binary covariates; more details can be found in Zeng et al. (2021); (iii) HSR: an observational study on the racial disparities of breast cancer screening, with 56,480 units and 17 binary covariates; more details can be found in Li et al. (2013). For each dataset, we fit a simple logistic model to estimate the propensity score $e(X)=\operatorname{\mathrm{e}xpit}(\beta_{0}+X^{\prime}\beta)$, and we provide the density plot and Q-Q plot in Figure 5. Normality holds approximately in all the cases.
Figure 5: Density and Q-Q plots of the fitted logit propensity scores, a linear combination of covariates. The proximity of points to the reference line in the Q-Q plot shows the closeness of the distribution to a normal distribution.